• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Markowitz

    Log entry time 08:00:17 on October22,2005

    Entry number 157671

    keyword=owl shift summary

    At start of shift, left septum is off. MCC cannot put beam into hall. MCC spent 4 hours tuning beam into hall, required adjustments due to compton. We took out the target to make tuning easier. Finally we turned off both septum, put beam onto dump, turned on raster, put back in target and started running in single arm.

    06:15 Q3 tripped, stopped run 5809.

    An access was taken to reboot Q3; a second access was also needed.

    07:15 Power permit, asked to put IMWP in. 07:38 started run 5810.

    07:45 Q3 tripped again, stopped run 5810.

    At end of shift: right HRS Q3 is tripped, left septum is still not cooling, Ed Folts is on his way in, and we called for the controlled access.